# Surprise-Adaptive Intrinsic Motivation for Unsupervised Reinforcement Learning

### v1

@@ 5.2 Adaptive Entropy Control
Our results have shown that the S-Adapt agent can successfully recreate the performance of the S-Min and the S-Max agents in their respective didactic environments. Next, we investigate controlling entropy across the MinAtar benchmark, shown in Figure 5. Notably, these environments were not constructed with any particular entropy regime in mind. Thus, these results provide an estimate of how the proposed algorithm could perform in an arbitrarily chosen environment.
@@
Figure 5: Average episode return (left) and surprise (right) versus environment interactions (average over 5 seeds, with one shaded standard deviation) in the MinAtar suite of environments. The S-Adapt agent indeed demonstrates emergent behaviors in certain environments, such as Freeway where it achieves rewards on par with that of the Extrinsic agent. However, in certain environments, like Seaquest, Space Invaders and Asterix, the extrinsic reward is not closely correlated with entropy control, with the Random agent and the Extrinsic agent achieving similar entropy at the end of training
Here again, we see that the S-Adapt agent can select the objective with the greatest controllable entropy. Though the difference between S-Min and S-Max in terms of divergence with the Random agent is not as strong in some environments, the S-Adapt agent consistently chooses the objective with the relatively larger change in entropy.
Next, we investigate the ability of the S-Adapt agent to control entropy given image-based observations as shown in Figure 6. The results show that the S-Adapt agent demonstrates emergent behavior and achieves slightly higher returns than the S-Max agent, which shows the effectiveness of our method given high dimensional image observations which are difficult to encode as binary observations, moreover, the results shows that the S-Adapt agent learns useful behaviors from a Gaussian distribution that is used for modeling the state marginal.
@@ 5.3 Emergent Behaviour
Finally, for these objectives to be useful, it is important that they correlate with the emergence of interesting behaviors. Indeed, we note that the extrinsic rewards in the didactic environments generally correlate closely with one of two single-objective agents. This suggests that these environments have good potential for entropy-based control to elicit emergent behaviors. Importantly, however, the extrinsic reward does not correlate well with either S-Min or S-Max in all environments. In Maze, S-Max achieves high rewards, while in Butterflies and Tetris, S-Min achieves high rewards. On the other hand, the S-Adapt agent achieves high task rewards, on par or better than the Extrinsic agent across all didactic environments.
Additionally, in some MinAtar environments, the entropy-based agents exhibit emergent behavior similar to that of the Extrinsic agent. In the Freeway environment, the S-Adapt agent achieves competitive rewards with the Extrinsic agent. A similar result is observed in Breakout. However, other environments, like Space Invaders and Seaquest do not appear to be good candidates for intrinsic entropy control, since the Extrinsic agent achieves similar entropy to that of the Random agent.
In Atari Freeway, the emergent behavior of the S-Adapt agent is qualitatively different from both S-Max and S-Min agents, the S-Adapt agent solves the game more frequently than the S-Max agent, this hints that mixing entropy maximization and minimization in one adaptive objective induces emergent behaviors which can not be learned by exclusively optimizing for surprise minimization or maximization.
@@ 6 Conclusion
Our experiments demonstrate encouraging results for a surprise-adaptive agent. The S-Adapt agent can select the objective with the more controllable landscape across both didactic environments and MinAtar environments. Moreover, the S-Adapt agent inherits the emergent behaviors of the single-objective agents, achieving high rewards across all didactic environments, which neither of the single-objective agents nor the extrinsic agent is able to achieve. Further work is needed to understand exactly under what conditions such emergent behaviors can manifest, and how to elicit them more reliably in arbitrary environments like MinAtar. Moreover, an interesting extension to this work would be to apply an adaptive agent in the continual learning setting, where adaptation can occur at any time, not only at episode end.
@@ A.1 Training Details
All agents were trained using DQN (Mnih et al., 2015). Reward values are normalized by subtracting the rolling mean and dividing by the standard deviation before fitting the Q network. For the S-Adapt agent, we use the original UCB algorithm with exploration coefficient 2 in maze (large) and MinAtar environments, for all other environments we set the exploration coefficien to $\sqrt{(}2)$. we trained all agents using the implementation from CleanRL Huang et al. (2022). We trained all agents with a learning of 0.0001, a discount factor of 0.99, a batch size of 32, a replay buffer size of 1M, and for 10M environment interactions
@@ A.2 Environments
We take the Tetris environment directly from the implementation provided by the authors of (Berseth et al., 2021). In this environment, the agent receives 0 at all steps, except for a losing step which results in a -100 reward. The maximum episode length is 200. Environment observations and the sufficient statistic of the state marginal are flattened before being fed into two independent two-layer MLP with hidden dimensions 120 and 84, the outputs of the MLPs are concatenated and passed through a linear layer that outputs the Q-value.
@@
The size of the small maze is 10x10 and the episode length is 100. Environment observations and the sufficient statistic of the state marginal are passed through two independent CNNs with a single convolutional layer, the outputs of the CNNs are concatenated and passed through a single-layer MLP with hidden dimension 512 that outputs the Q-value.
The size of the large maze is 32x32 and the episode length is 250. Environment observations and the sufficient statistic of the state marginal are passed through two independent CNNs with three convolutional layers, the outputs of the CNNs are concatenated and passed through a single-layer MLP with hidden dimension 512 that outputs the q-value.
@@
We constructed the custom Butterflies environment (small and large) using the Griddly platform (Bamford, 2021). A pixel-rendering of the small and large mazes used in our experiments can be found in Figure 7. The task reward in both environments is +1 when the agent catches a butterfly and 0 otherwise. The size of the small map is 10x10 and the episode length is 100. Environment observations and the sufficient statistic of the state marginal are passed through two independent CNNs with a single convolutional layer, the outputs of the CNNs are concatenated and passed through a single-layer MLP with hidden dimension 512 that outputs the Q-value. The size of the large map is 32x32 and the episode length is 500. Environment observations and the sufficient statistic of the state marginal are passed through two independent CNNs with three convolutional layers, the outputs of the CNNs are concatenated and passed through a single-layer MLP with hidden dimension 512 that outputs the q-value.
@@ A.3 Estimation of State Marginal Distribution
In all environments, the observed state $s_{t}$ is a binary entity map of size $H\times W\times C$, where $H$ is the height of the map, $W$ is the width of the map and $C$ is the number of channels, with each channel representing a single object type in the environment. A value of one is set in the $(h,w)$ position of channel $c$ (denoted $s_{t}^{h,w,c}$) if an object of type $c$ currently occupies the $(h,w)$ position in the map, and zero otherwise. The state marginal distribution is estimated as $H\times W\times C$ independent Bernoulli distributions, with probability $p_{t}^{h,w,c}=\frac{\sum_{t^{\prime}=0}^{t}s^{h,w,c}_{t^{\prime}}}{t}$, which constitutes a sufficient statistic for the Bernoulli distribution. Hence, the sufficient statistic of the entire state marginal distribution is given by $\theta_{t}=\{p_{t}^{h,w,c}:h\in H,w\in W,c\in C\}$ and is the same shape as the observations $s_{t}$.
The choice of the Bernoulli distribution is justified by the binary nature of the data. However, we perform an ablation using a Gaussian distribution as an alternative to confirm the validity of this choice (Figure 8)
@@
Appendix B Butterflies Density
@@
Figure 9: Average episode return of the S-Adapt agent (average over 5 seeds, with one shaded standard deviation) over various butterflies densities in the Butterflies (large) environment. At (very) high density (Figures 9(a) and 9(b)), the Random agent resembles the S-Min agent and catches large number of butterflies as indicated by the high episode return. Hence, the S-Adapt agent converges to surprise-maximization to induce large absolute difference in entropy from the random agent and avoids butterflies as indicated by the low episodic return. In contrast, at very low density (Figure 9(e)), the random agent is unable to catch butterflies and resembles the S-Max agent. The S-Adapt agent converges to surprise-minimization and almost catches all the butterflies as indicated by the high episodic return, at medium and low densities (Figures 9(c) and 9(d)), the S-Adapt agent oscillates between surprise-maximization and surprise-minimization as they roughly induce the same absolute difference in entropy.

### v2

@@ 5.2 Adaptive Entropy Control
Our results have shown that the S-Adapt agent can successfully recreate the performance of the S-Min and the S-Max agents in their respective didactic environments. Next, we investigate controlling entropy across the MinAtar benchmark, shown in Figure 5. Notably, these environments were not constructed with any particular entropy regime in mind. Thus, these results are demonstrative of how the proposed algorithm could perform in an arbitrarily chosen environment.
@@
Figure 5: Average episode return (left) and surprise (right) versus environment interactions (average over 5 seeds, with one shaded standard deviation) in the MinAtar suite of environments. In all environments the S-Adapt agent is able to select the direction for entropy optimization which is most controllable, as demonstrated by the change in entropy from the beginning to the end of training. The S-Adapt agent indeed demonstrates emergent behaviors in certain environments, such as Freeway where it achieves rewards on par with that of the Extrinsic agent. However, in certain environments, like Seaquest, Space Invaders and Asterix, the extrinsic reward is not closely correlated with entropy control, with the Random agent and the Extrinsic agent achieving similar entropy.
Here again, we see that the S-Adapt agent can reliably select the objective with the greatest controllable entropy. Though the difference between S-Min and S-Max agents in terms of divergence with the Random agent is not as strong in some environments, the S-Adapt agent consistently chooses the objective with the relatively larger change in entropy. This provides confirmation that our bandit algorithm can successfully select for controllable entropy in arbitrary environments.
@@ 5.3 Emergent Behaviour
Finally, for these objectives to be useful, it is important that they correlate with the emergence of interesting behaviors. Indeed, we note that the extrinsic rewards in the didactic environments generally correlate closely with one of two single-objective agents (Figures 2, 3 and 4). This suggests that these environments have good potential for entropy-based control to elicit emergent behaviors. Importantly, however, the extrinsic reward does not correlate well with strictly one of S-Min or S-Max in all environments. In Maze, S-Max achieves high rewards, while in Butterflies and Tetris, S-Min achieves high rewards. On the other hand, the S-Adapt agent achieves high task rewards, on par or better than the Extrinsic agent across all didactic environments.
Additionally, in some MinAtar environments, the entropy-based agents exhibit emergent behavior similar to that of the Extrinsic agent. In the Freeway environment (Figure 5(a)), the S-Adapt agent achieves competitive rewards with the Extrinsic agent. A similar result is observed in Breakout (Figure 5(d)). However, other environments, like Space Invaders and Seaquest (Figures 5(c) and 5(b)) do not appear to be good candidates for intrinsic entropy control, since the Extrinsic and Random agents achieve similar entropy.
Finally, we investigate the emergence of interesting behaviors in a more complex, image-based environment using Atari Freeway (Figure 6) as a case study. Unlike the previous environments, observations in pixel space are non-binary and hence cannot be modeled using Bernoulli distributions. Instead, we model the state marginal using a Gaussian distribution (see Section A.3 for more details). The results show that both the S-Max and S-Adapt agents achieve respectable results as compared to the extrinsic agent. Moreover, in this environment, the emergent behavior of the S-Adapt agent is qualitatively different from both S-Max and S-Min agents; The S-Adapt agent solves the game more frequently than the S-Max agent. This hints that mixing entropy maximization and minimization in one adaptive objective induces emergent behaviors that cannot be learned by exclusively optimizing for surprise minimization or maximization alone.
@@ 6 Conclusion
Our experiments demonstrate encouraging results for a surprise-adaptive agent. The S-Adapt agent can select the objective with the more controllable landscape across both didactic environments and benchmark environments. Moreover, the S-Adapt agent inherits the emergent behaviors of the single-objective agents, achieving high rewards across all didactic environments, which neither of the single-objective agents nor the extrinsic agent is able to achieve. Further work is needed to understand exactly under what conditions such emergent behaviors can manifest, and how to elicit them more reliably in arbitrary environments like MinAtar. Possible directions for improvement here could include better methods for estimating the state marginal distribution with more accuracy. Moreover, an interesting extension to this work would be to apply an adaptive agent in the continual learning setting, where adaptation can occur at any time, not only at episode end.
Acknowledgments
We want to acknowledge funding support from NSERC, FRQNT, and CIFAR and compute support from the Digital Research Alliance of Canada, Mila IDT and NVidia.
@@ A.1 Training Details
All agents were trained using DQN (Mnih et al., 2015). Reward values are normalized by subtracting the rolling mean and dividing by the standard deviation before fitting the Q network. For the S-Adapt agent, we use the original UCB algorithm with exploration coefficient 2 in the Maze (large) and MinAtar environments, for all other environments we set the exploration coefficient to $\sqrt{2}$. we trained all agents using the implementation of DQN from CleanRL (Huang et al., 2022). We trained all agents with a learning rate of 0.0001 with Adam optimizer, a discount factor of 0.99, a batch size of 32, a replay buffer size of 1M, and for 10M environment interactions. We use epsilon-greedy for exploration with a linearly decaying epsilon from a value of 1 to 0.01, decaying over the first 10% of timesteps in all environments except MinAtar and Atari which decays over the first 50% of time steps. Model architecture details for each environment are provided in the next section.
@@ A.2 Environments
We take the Tetris environment directly from the implementation provided by the authors of (Berseth et al., 2021). In this environment, the agent receives 0 at all steps, except for a losing step which results in a -100 reward. The maximum episode length is 200. Environment observations and the sufficient statistic of the state marginal are flattened before being fed into two independent two-layer MLPs with hidden dimensions 120 and 84. The outputs of the MLPs are concatenated and passed through a linear layer that outputs the Q-value.
@@
The size of the small maze is 10x10 and the episode length is 100. Environment observations and the sufficient statistic of the state marginal are passed through two independent CNNs with a single convolutional layer. The outputs of the CNNs are concatenated and passed through a single-layer MLP with hidden dimension 512 that outputs the Q-value.
The size of the large maze is 32x32 and the episode length is 250. Environment observations and the sufficient statistic of the state marginal are passed through two independent CNNs with three convolutional layers with kernel size of (3,3), a stride value of 2 and a padding value of 1. The outputs of the CNNs are concatenated and passed through a single-layer MLP with hidden dimension 512 that outputs the Q-value.
@@
We constructed the custom Butterflies environment (small and large) using the Griddly platform (Bamford, 2021). The task reward in both environments is +1 when the agent catches a butterfly and 0 otherwise.
The size of the small map is 10x10 and the episode length is 100, while the size of the large map is 32x32 and the episode length is 500. We use the same architecture as the Maze environment for estimating the Q-value.
MinAtar
In MinAtar environments, we use the same architecture as the Butterflies environments and we set the episode length to 500.
Atari
In Atari Freeway environment, we use the same architecture and pre-processing as in Mnih et al. (2015). We use the same multiple CNN architecture as the Maze environment for estimating the Q-values from the augmented state with sufficient statistics.
@@ A.3 Estimation of State Marginal Distribution
In all binary environments (Tetris, Maze, Butterflies, MinAtar), the observed state $s_{t}$ is a binary entity map of size $H\times W\times C$, where $H$ is the height of the map, $W$ is the width of the map and $C$ is the number of channels, with each channel representing a single object type in the environment. A value of one is set in the $(h,w)$ position of channel $c$ (denoted $s_{t}^{h,w,c}$) if an object of type $c$ currently occupies the $(h,w)$ position in the map, and zero otherwise. The state marginal distribution is estimated as $H\times W\times C$ independent Bernoulli distributions, with probability $p_{t}^{h,w,c}=\frac{\sum_{t^{\prime}=0}^{t}s^{h,w,c}_{t^{\prime}}}{t}$, which constitutes a sufficient statistic for the Bernoulli distribution. Hence, the sufficient statistic of the entire state marginal distribution is given by $\theta_{t}=\{p_{t}^{h,w,c}:h\in H,w\in W,c\in C\}$ and is the same shape as the observations $s_{t}$.
The choice of the Bernoulli distribution is justified by the binary nature of the data. However, we perform an ablation using a Gaussian distribution as an alternative to confirm the validity of this choice (Figure 8).
In the image-based environment (Atari Freeway), the observed state $s_{t}$ is an image. Here, we use a Gaussian distribution for the state marginal estimation. Using the same notation as above, the sufficient statistics for the Gaussian distribution are given by empirical mean and variance $\mu^{h,w,c}_{t}=\frac{\sum_{t^{\prime}=0}^{t}s_{t^{\prime}}^{h,w,c}}{t}$, $\sigma^{h,w,c}_{t}=\frac{\sum_{t^{\prime}=0}^{t}(\mu^{h,w,c}_{t}-s_{t^{\prime}}^{h,w,c})^{2}}{t}$. The sufficient statistic for the entire state marginal distribution is then given by $\theta_{t}=\{\mu_{t}^{h,w,c},\sigma_{t}^{h,w,c}:h\in H,w\in W,c\in C\}$.
@@
Appendix B Additional Experiments
Here we present additional results on the impact of butterfly density on the behavior of the S-Adapt agent.
@@ Appendix B Additional Experiments
Figure 9: Average episode return of the S-Adapt agent (average over 5 seeds, with one shaded standard deviation) over various butterflies densities in the Butterflies (large) environment. At (very) high density (Figures 9(a) and 9(b)), the Random agent resembles the S-Min agent and catches large number of butterflies as indicated by the high episode return. Hence, the S-Adapt agent converges to surprise-maximization to induce large absolute difference in entropy from the Random agent and avoids butterflies as indicated by the low episodic return. In contrast, at very low density (Figure 9(e)), the Random agent is unable to catch butterflies and resembles the S-Max agent. The S-Adapt agent converges to surprise-minimization and almost catches all the butterflies as indicated by the high episodic return. At medium and low densities (Figures 9(c) and 9(d)), the S-Adapt agent oscillates between surprise-maximization and surprise-minimization as they roughly induce the same absolute difference in entropy.